# Imperfect dark matter with higher derivatives

### v1

@@ Abstract
We introduce a higher-derivative action for dark matter whose energy–momentum tensor describes an imperfect fluid with nonzero pressure, energy flux, and anisotropic stress. In the limit where the higher-derivative couplings are switched off, the energy-momentum tensor reduces to pressureless dust. A systematic derivation follows from extending the singular conformal transformation used in the mimetic dark matter scenario to include higher-derivative terms while the resulting action is general and does not rely on the mimetic framework. On a homogeneous cosmological background, the dynamics coincides with that of pressureless dust, while in the presence of inhomogeneities the higher-derivative terms generate nonzero acceleration and vorticity, making it possible to avoid the formation of caustic singularities even if the strong energy condition satisfies. In particular, within the mimetic realization these terms resolve the usual caustic pathology of mimetic dark matter.
@@ 1 Introduction
The action for a perfect fluid in terms of hydrodynamical variables was first introduced by Schutz [4]. For the case of dust, we have [5, 6]
@@
where $g_{\mu\nu}$ is the spacetime metric and we work with metric signature (,+,+,+)(-,+,+,+). Taking variation with respect to the auxiliary field $\lambda$ or equivalently energy density $\rho$ gives the normalization condition for the four-velocity while taking variation with respect to the metric correctly reproduces the dust energy-momentum tensor (1.1) that is defined as $T^{\rm dust}_{\mu\nu}=-(2/\sqrt{-g})\delta{S}_{\rm dust}/\delta{g}^{\mu\nu}$.
@@
where $\tilde{g}_{\mu\nu}$ is an auxiliary metric. The inverse contravariant metric is given by $g^{\mu\nu}=-\left(\tilde{g}^{\alpha\beta}u_{\alpha}u_{\beta}\right)^{-1}\tilde{g}^{\mu\nu}$ which after contracting both sides with $u_{\mu}{u}_{\nu}$ gives $u^{\alpha}u_{\alpha}=-1$. Note that we cannot express auxiliary metric $\tilde{g}_{\mu\nu}$ in terms of the physical metric $g_{\mu\nu}$. This shows that the conformal transformation (1.3) is singular and, therefore, imposing it at the level of the action yields a dynamically inequivalent theory. We discuss this in detail in the next section. Substituting (1.3) directly into the Einstein–Hilbert action [^1] We work in units $\hbar=c=1$ such that $M_{\rm Pl}=1/\sqrt{8\pi{G}}$ is the reduced Planck mass. $S_{\rm E.H.}=\tfrac{M_{\rm Pl}^{2}}{2}\int\sqrt{-g}R$ yields field equations equivalent to general relativity sourced by pressureless dust; the dust arises from the isolated conformal degree of freedom of the metric. Subsequently, it was shown that applying the singular transformation (1.3) to the Einstein–Hilbert action is classically equivalent to working entirely with the physical metric and supplementing the Einstein–Hilbert action with the dust action (1.2) [9].
@@
Because congruences of geodesics generally focus, dust flows develop caustics (shell crossing) in finite time. When the dust is identified with the conformal mode, such caustics manifest as singular behavior of the mimetic field, signaling a breakdown of the mimetic reconstruction of the metric [10]. To prevent caustic formation, non-vanishing acceleration (deviation from the geodesics equation (1.4)) or non-vanishing vorticity is necessary. For example, one may add $F_{\mu\nu}=2\partial_{[\mu}u_{\nu]}$ to give dynamics to the transverse modes leading to non-vanishing vorticity which might avoid formation of caustics [8, 11, 12]. [^2] One may also consider gauge–field extensions of the mimetic construction, which are generally free of caustics pathology [13, 14, 15, 16, 17]. However, in those cases the mimetic sector no longer reproduces dark matter phenomenology. In this paper, we instead focus on the role of higher-derivative couplings. In order to do so, we first systematically take into account higher-derivative terms within the mimetic framework, and then show that higher-derivative terms generally induce both acceleration and vorticity, thereby regulating the flow allowing caustics to be avoided.
@@ 2 Singular disformal transformations
To find the singular limit, we look at the eigenvalue problem [23, 24]
@@
$J^{\rho\sigma}_{\mu\nu}=A\,\delta^{(\rho}_{\mu}\delta^{\sigma)}_{\nu}-\left[A_{\tilde{X}}\tilde{\phi}^{\rho}\tilde{\phi}^{\sigma}+2A_{\tilde{Y}}\tilde{\phi}^{(\rho}\tilde{X}^{\sigma)}+A_{\tilde{Z}}\tilde{X}^{\rho}\tilde{X}^{\sigma}+2A_{\tilde{Z}}\tilde{X}^{\alpha}\tilde{\nabla}_{\alpha}\left(\tilde{\phi}^{\rho}\tilde{\phi}^{\sigma}\right)\right]\tilde{g}_{\mu\nu}\,.$ (2.15)
@@
We now rewrite the result (2.17) in terms of quantities built from the physical metric. To leading order in a derivative expansion, one finds
@@
Using these relations, we obtain
@@ 3 Higher-derivative mimetic gravity
$\begin{split}C_{\phi\phi}\equiv\left(Xf\right)_{,X}-\frac{1}{\lambda}\nabla_{\alpha}\left(\lambda C_{\phi{X}}\phi^{\alpha}+2\lambda C_{XX}X^{\alpha}\right)\,,\quad C_{\phi{X}}\equiv\left(Xf\right)_{,Y}\,,\quad C_{XX}\equiv\left(Xf\right)_{,Z}\,.\end{split}$
@@
The action (3.1) contains up to second derivatives of $\phi$ (through $X_{\mu}$ and $Z$) and the corresponding equations of motion involve up to fourth-order derivatives of $\phi$. Such higher-order equations typically signal an Ostrogradsky instability [34]. However, the situation changes in the mimetic scenario since the system is degenerate [11, 35, 36, 37, 38, 39, 31, 40]. In our construction the higher-derivative dependence appears only through the constraint $X\,f(Y/X^{2},Z/X^{3})=1$ enforced by the equation of the auxiliary field $\lambda$. This imposes a strong (algebraic) relation among $X$, $Y=\!\phi^{\alpha}X_{\alpha}$, and $Z=\!X^{\alpha}X_{\alpha}$, i.e. among second derivatives of $\phi$ and its first derivatives. As a result, after using the constraint $Xf=1$ in the field equations, all higher-derivative contributions can be reduced to the lower-derivative quantities, so that on shell no Ostrogradsky-like ghost propagates. This is most transparent in unitary gauge, $\phi(x,t)=t$. Then $X=g^{00}=-N^{-2}$, where $N$ is the lapse function, while $Y$ and $Z$ depend on the time and spatial derivatives of $N$. The constraint $Xf(Y/X^{2},Z/X^{3})=1$ becomes an algebraic relation relating $\dot{N}$ to $N$ and its spatial derivatives, thereby eliminating time derivatives of the lapse from the dynamics. A complete Hamiltonian analysis would provide a rigorous confirmation of the constraint structure for (3.1), and we leave this for future work.
@@
Moreover, the fact that, in our construction, all higher-derivative effects enter exclusively through the modified constraint term in the action (3.1) makes it different from other approaches that add higher-derivative operators directly to the Lagrangian while keeping the original mimetic constraint unchanged [41, 19, 42, 20, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 36, 53, 54, 55, 56]. The theory (3.1) is therefore qualitatively different from those earlier proposals. As an illustration, if the mimetic constraint is left unchanged so that $\phi^{\alpha}\phi_{\alpha}=-1$, then $X_{\mu}=2\phi^{\alpha}\phi_{\alpha\mu}=0$, and the energy–momentum tensor reduces to that of dust with $\rho=2\lambda$ and $u^{\mu}=\phi^{\mu}$. In the usual mimetic case $u_{\mu}=\partial_{\mu}\phi$ is hypersurface-orthogonal (hence irrotational), and the flow is geodesic, so caustics generically form. By contrast, in our model (3.1) the constraint is modified; $u^{\mu}$ is no longer aligned with $\partial^{\mu}\phi$ and need not be hypersurface-orthogonal, so the higher-derivative terms generically induce acceleration and vorticity, allowing caustic formation to be avoided. We will give detail analysis about caustics in Sec. 5.
4 $3+1$ decomposition of energy-momentum tensor
@@
where $\rho=u^{\alpha}{u}^{\beta}{T}_{\alpha\beta}$ can be found for a given observer with four-velocity $u^{\mu}$. In this respect, our setup provide dark matter in cosmological homogeneous and isotropic background. By contrast, for standard dust (1.1) one has $p=0$, $q^{\mu}=0$, and $\pi^{\mu\nu}=0$ identically, independent of the background or perturbations. In our setup these quantities vanish only on the homogeneous background and are generically nonzero once inhomogeneities are present (e.g., at linear order in perturbations). Thus the model mimics pressureless dark matter at the background level while allowing nontrivial perturbations.
@@ 4.1 Comoving frame
$\varphi_{\perp}=-\sqrt{-X}\,,\qquad\varphi^{\mu}=0\,,\qquad\chi_{\perp}=\tfrac{Y}{\sqrt{-X}}\,,\qquad\chi^{\mu}=X^{\mu}-\tfrac{Y}{X}\phi^{\mu}\,$, (4.13)
@@
$\rho$ $=2\lambda\!\left(C_{\phi\phi}X+C_{\phi X}Y+C_{XX}\tfrac{Y^{2}}{X}\right)\,$,
$q^{\mu}$ $=-\lambda\sqrt{-X}\left(C_{\phi X}+2C_{XX}\tfrac{Y}{X}\right)\chi^{\mu},$
$\tau^{\mu\nu}$ $=-2\lambda\,C_{XX}\,\chi^{\mu}\chi^{\nu},\qquad p=-\tfrac{2\lambda}{3}C_{XX}\,\chi_{\alpha}\chi^{\alpha},$
$\pi^{\mu\nu}$ $=-2\lambda\,C_{XX}\!\left(\chi^{\mu}\chi^{\nu}-\tfrac{1}{3}\chi_{\alpha}\chi^{\alpha}h^{\mu\nu}\right).$
This frame makes the higher-derivative effects manifest: since $\varphi^{\mu}=0$, any nonzero pressure, energy flux, and anisotropic stress arise entirely from $\chi^{\mu}$, the spatial projection of $X^{\mu}$. Note that $X_{\mu}=2\,\phi^{\alpha}\phi_{\alpha\mu}$ and $\chi^{\mu}$ characterizes the spatial second derivatives of the scalar field.

### v2

@@ Abstract
We introduce a higher-derivative action for dark matter whose energy-momentum tensor describes an imperfect fluid with nonzero pressure, energy flux, and anisotropic stress. In the limit where the higher-derivative couplings are switched off, the energy-momentum tensor reduces to pressureless dust. A systematic derivation follows from extending the singular conformal transformation used in the mimetic dark matter scenario to include higher-derivative terms while the resulting action is general and does not rely on the mimetic framework. On a homogeneous cosmological background, the dynamics coincides with that of pressureless dust, while in the presence of inhomogeneities the higher-derivative terms generate nonzero acceleration and vorticity, making it possible to avoid the formation of caustic singularities even if the strong energy condition satisfies. In particular, within the mimetic realization these terms can resolve the usual caustic pathology of mimetic dark matter.
@@ 1 Introduction
The action for a perfect fluid in terms of hydrodynamical variables was first introduced by Schutz [4]. For the case of irrotational dust, we have [5, 6]
@@
where $g_{\mu\nu}$ is the spacetime metric, $u_{\mu}$ is a hypersurface-orthogonal four-velocity, and we work with metric signature (,+,+,+)(-,+,+,+). Taking variation with respect to the auxiliary field $\lambda$ or equivalently energy density $\rho$ gives the normalization condition for the four-velocity while taking variation with respect to the metric correctly reproduces the dust energy-momentum tensor (1.1) that is defined as $T^{\rm dust}_{\mu\nu}=-(2/\sqrt{-g})\delta{S}_{\rm dust}/\delta{g}^{\mu\nu}$.
@@
where $\tilde{g}_{\mu\nu}$ is an auxiliary metric. The inverse contravariant metric is given by $g^{\mu\nu}=-\left(\tilde{g}^{\alpha\beta}u_{\alpha}u_{\beta}\right)^{-1}\tilde{g}^{\mu\nu}$ which after contracting both sides with $u_{\mu}{u}_{\nu}$ gives $u^{\alpha}u_{\alpha}=-1$. Note that we cannot express auxiliary metric $\tilde{g}_{\mu\nu}$ in terms of the physical metric $g_{\mu\nu}$. This shows that the conformal transformation (1.3) is singular and, therefore, imposing it at the level of the action yields a dynamically inequivalent theory. We discuss this in detail in the next section. Substituting (1.3) directly into the Einstein-Hilbert action [^1] We work in units $\hbar=c=1$ such that $M_{\rm Pl}=1/\sqrt{8\pi{G}}$ is the reduced Planck mass. $S_{\rm E.H.}=\tfrac{M_{\rm Pl}^{2}}{2}\int{\rm d}^{4}x\sqrt{-g}R$ yields field equations equivalent to general relativity sourced by pressureless dust; the dust arises from the isolated conformal degree of freedom of the metric. Subsequently, it was shown that applying the singular transformation (1.3) to the Einstein-Hilbert action is classically equivalent to working entirely with the physical metric and supplementing the Einstein-Hilbert action with the dust action (1.2) [9].
@@
Because congruences of geodesics generally focus, dust flows develop caustics (shell crossing) in finite time. When the dust is identified with the conformal mode, such caustics manifest as singular behavior of the mimetic field, signaling a breakdown of the mimetic reconstruction of the metric [10]. To prevent caustic formation, non-vanishing acceleration (deviation from the geodesics equation (1.4)) or non-vanishing vorticity is necessary. For example, one may add $F_{\mu\nu}=2\partial_{[\mu}u_{\nu]}$ to give dynamics to the transverse modes leading to non-vanishing vorticity which might avoid formation of caustics [8, 11, 12]. [^2] One may also consider gauge-field extensions of the mimetic construction, which are generally free of caustics pathology [13, 14, 15, 16, 17]. However, in those cases the mimetic sector no longer reproduces dark matter phenomenology. In this paper, we instead focus on the role of higher-derivative couplings. In order to do so, we first systematically take into account higher-derivative terms within the mimetic framework, and then show that higher-derivative terms generally induce both acceleration and vorticity, thereby regulating the flow allowing caustics to be avoided.
@@ 2 Singular disformal transformations
To find the singular limit, we look at the eigenvalue problem [23, 24, 25]
@@
$J^{\rho\sigma}_{\mu\nu}=A\,\delta^{(\rho}_{\mu}\delta^{\sigma)}_{\nu}-\left[A_{\tilde{X}}\tilde{\phi}^{\rho}\tilde{\phi}^{\sigma}+2A_{\tilde{Y}}\tilde{\phi}^{(\rho}\tilde{X}^{\sigma)}+A_{\tilde{Z}}\tilde{X}^{\rho}\tilde{X}^{\sigma}+2A_{\tilde{Z}}\tilde{X}^{\alpha}\tilde{\nabla}_{\alpha}\left(\tilde{\phi}^{\rho}\tilde{\phi}^{\sigma}\right)\right]\tilde{g}_{\mu\nu}+\cdots\,$, (2.15)
where the ellipses denote contributions involving the third-order terms listed in Eq. (2.14).
@@
To leading order in a derivative expansion, one finds
@@
Note that, similar to the usual mimetic transformation (1.3), the generalized higher-derivative conformal transformation $g_{\mu\nu}=A\,\tilde{g}_{\mu\nu}$ with $A$ given by (2.17) is invariant, at the order retained in this section, under a local Weyl rescaling of the auxiliary metric $\tilde{g}_{\mu\nu}\to\Omega^{2}(x)\tilde{g}_{\mu\nu}$: indeed, $\tilde{X}\to\Omega^{-2}\tilde{X}$ while the dimensionless ratios $\tilde{Y}/\tilde{X}^{2}$ and $\tilde{Z}/\tilde{X}^{3}$ are invariant at this order, so that $A\to\Omega^{-2}A$ and consequently $g_{\mu\nu}=A\,\tilde{g}_{\mu\nu}$ remains unchanged. This auxiliary Weyl redundancy is the hallmark of the non-invertibility of the map within our approximation and, as in standard mimetic gravity, it is precisely what isolates the conformal mode.
We now rewrite the result (2.17) in terms of quantities built from the physical metric. Using (2.18) in (2.17), we obtain
@@ 3 Higher-derivative mimetic gravity
$\begin{split}C_{\phi\phi}\equiv\left(Xf\right)_{,X}-\frac{1}{\lambda}\nabla_{\alpha}\left(\lambda C_{\phi{X}}\phi^{\alpha}+2\lambda C_{XX}X^{\alpha}\right)\,,\quad C_{\phi{X}}\equiv\left(Xf\right)_{,Y}\,,\quad C_{XX}\equiv\left(Xf\right)_{,Z}\,.\end{split}$ (3.5)
@@
The action (3.1) contains up to second derivatives of $\phi$ (through $X_{\mu}$ and $Z$) and the corresponding equations of motion involve up to fourth-order derivatives of $\phi$. Such higher-order equations typically signal an Ostrogradsky instability [35]. However, the situation changes in the mimetic scenario since the system is degenerate [11, 36, 37, 38, 39, 40, 32, 41]. In our construction the higher-derivative dependence appears only through the constraint $X\,f(Y/X^{2},Z/X^{3})=1$ enforced by the equation of the auxiliary field $\lambda$. This imposes a strong (algebraic) relation among $X$, $Y=\!\phi^{\alpha}X_{\alpha}$, and $Z=\!X^{\alpha}X_{\alpha}$, i.e. among second derivatives of $\phi$ and its first derivatives. As a result, after using the constraint $Xf=1$ in the field equations, all higher-derivative contributions can be reduced to the lower-derivative quantities. This might (or might not) be enough to eliminate the would-be Ostrogradsky ghost. Therefore, a complete Hamiltonian analysis would provide a rigorous confirmation of the constraint structure for (3.1), and we leave this for future work.
On the other hand, there is a particularly simple subclass of the theory in which the potentially dangerous higher-time-derivative structure is absent, so that the would-be Ostrogradsky mode is eliminated already at the level of the operator basis. This is obtained by restricting the constraint function to depend only on the following combination
${\cal C}\;\equiv\;\frac{Z}{X^{3}}-\left(\frac{Y}{X^{2}}\right)^{2}\;.$ (3.7)
Introducing the transverse spatial vector
$V_{\mu}\;\equiv\;X_{\mu}-\frac{Y}{X}\phi_{\mu},\qquad\phi^{\mu}V_{\mu}=0,$ (3.8)
one immediately finds
${\cal C}=\frac{V_{\mu}V^{\mu}}{X^{3}}\,.$ (3.9)
In this form, it is clear that $\cal C$} is constructed solely from the component of $X_{\mu}$ orthogonal to $\phi_{\mu}$. In particular, it does not probe the “longitudinal” component along $\phi_{\mu}$, which is responsible for higher time-derivative contributions. This is especially transparent in unitary gauge, $\phi=t$. Using the ADM decomposition ${{\rm d}}s^{2}=-N^{2}{{\rm d}}t^{2}+h_{ij}({{\rm d}}x^{i}+N^{i}{{\rm d}}t)({{\rm d}}x^{j}+N^{j}{{\rm d}}t)$, one finds $X=-1/N^{2}$ and ${\cal C}=h^{ij}\partial_{i}X\partial_{j}X/X^{3}$, which involves no time derivative of the lapse (and hence no second time derivatives of the scalar field). The situation is therefore closely analogous to the standard mimetic case: the mimetic constraint reduces to
$Xf\left({\cal C}\right)=1\,$, (3.10)
and, in unitary gauge, it is a first-order equation for the scalar field. Accordingly, the number of propagating degrees of freedom in unitary gauge matches the usual mimetic scenario: two tensor modes and one scalar mode.
It is worth noting that the combination (3.7) is not contained in Horndeski or, more generally, in Degenerate Higher-Order Scalar-Tensor (DHOST) theories [42], but instead arises in the context of U-DHOST theories [43, 44, 45]. In U-DHOST, the theory propagates three degrees of freedom (two tensors and one scalar) in unitary gauge, while beyond unitary gauge an additional mode appears; however, this extra mode is instantaneous, with an elliptic equation of motion, and is therefore harmless. Moreover, the “scordatura” mechanism, which provides higher spatial derivatives while keeping higher time derivatives suppressed below a cutoff [46, 47, 48], is built in automatically in U-DHOST. Consequently, the subset of our higher-derivative setup restricted to (3.7) can also yield healthy higher-spatial-derivative behavior in the high-energy regime where higher-derivative operators become relevant.
The covariant equations of motion significantly simplify once we restrict our setup to $f({\cal C})$: the coefficients in (3.5) simplify to
$C_{\phi X}=-\frac{2Y}{X}\,C_{XX}=-\frac{2Y}{X^{3}}\,f_{\cal C}\,$, (3.11)
and the energy-momentum tensor (3.4) can be written in the compact form
$T_{\mu\nu}=-2\lambda\Big[\widehat{C}_{\phi\phi}\,\phi_{\mu}\phi_{\nu}+\frac{f_{\cal C}}{X^{2}}\,V_{\mu}V_{\nu}\Big],$ (3.12)
where
$\widehat{C}_{\phi\phi}\equiv C_{\phi\phi}-\frac{Y^{2}}{X^{2}}C_{XX}=f-3{\cal C}f_{\cal C}-\frac{2}{\lambda}\nabla_{\alpha}\!\left(\lambda\,\frac{f_{\cal C}}{X^{2}}\,V^{\alpha}\right).$ (3.13)
The result (3.12) shows that $\phi_{\mu}$ is the unique time-like direction and can therefore be used as the fluid four-velocity, $u_{\mu}\propto\phi_{\mu}$. The remaining vector $V_{\mu}$ is purely spatial and thus provides the only additional direction available to construct imperfect contributions, which are entirely encoded in the single tensor structure $V_{\mu}V_{\nu}$. We will discuss the associated fluid interpretation and its consequences in more detail in the next section.
Moreover, the fact that in our construction all higher-derivative effects enter exclusively through the (modified) constraint sector in the action (3.1) makes it qualitatively different from the standard “higher-derivative mimetic” setups, where one keeps the original mimetic constraint X=1X=-1 (implemented via a term $\lambda(X+1)$) and adds higher-derivative operators directly in the Lagrangian, e.g. through functions of $\Box\phi$ and/or other higher-derivative combinations [49, 19, 50, 20, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 37, 61, 62, 63, 64]. In those approaches the constraint is fixed to $X\equiv\phi_{\mu}\phi^{\mu}=-1$, so that $\nabla_{\mu}X=2\phi^{\alpha}\phi_{\alpha\mu}=0$ identically. In particular, for the commonly used higher-derivative scalars $Y\equiv 2\phi^{\alpha}\phi^{\beta}\phi_{\alpha\beta}$ and $Z\equiv 4\phi^{\alpha}\phi^{\beta}\phi_{\alpha\eta}\phi_{\beta}{}^{\eta}$ one has $Y=0$ and $Z=0$ on the constraint surface, and similarly the “longitudinal” building blocks $\vartheta_{n}\equiv\phi^{\alpha}\phi^{\beta}(\phi_{\alpha\beta}^{n})$ vanish for $n\geq 1$. This observation was emphasized in the DHOST extension of mimetic gravity [62], where, precisely because $\vartheta_{n}=0$ when X=1X=-1, the nontrivial higher-derivative operators may be chosen to depend instead on the trace-type combinations $\chi_{n}\equiv g^{\alpha\beta}(\phi_{\alpha\beta}^{n})$ (in particular $\chi_{1}=\Box\phi$). By contrast, in our theory the constraint itself is deformed (schematically of the form $Xf=1$), so $X$ is not fixed to a constant and $\nabla_{\mu}X$ need not vanish. As a result, $Y$ and $Z$ (equivalently $\vartheta_{1}$ and $\vartheta_{2}$) are generically nonzero and become natural, non-redundant carriers of higher-derivative effects; this is in fact what one finds by a systematic analysis of the relevant singular limit, which indicates that the deformation of the constraint is essential and that the appropriate higher-derivative information is encoded in the “longitudinal” structures rather than being pushed into $\Box\phi$-type terms.
@@
As an illustration of the physical distinction, if the constraint is left unchanged so that $\phi^{\alpha}\phi_{\alpha}=-1$, then $X_{\mu}=2\phi^{\alpha}\phi_{\alpha\mu}=0$ and the energy-momentum tensor reduces to that of dust with $\rho=2\lambda$ and $u^{\mu}=\phi^{\mu}$. In the usual mimetic case $u_{\mu}=\partial_{\mu}\phi$ is hypersurface-orthogonal (hence irrotational) and the flow is geodesic, so caustics generically form. By contrast, in our model (3.1) the constraint is modified; $u^{\mu}$ is no longer aligned with $\partial^{\mu}\phi$ and need not be hypersurface-orthogonal, so the higher-derivative contributions generically induce acceleration and vorticity, allowing caustic formation to be avoided. A detailed analysis of caustics will be presented in Sec. 5.
4 $1+3$ decomposition of energy-momentum tensor
@@
where $\rho=u^{\alpha}{u}^{\beta}{T}_{\alpha\beta}$ can be found for a given observer with four-velocity $u^{\mu}$. In this respect, our setup provide dark matter in cosmological homogeneous and isotropic background. By contrast, for standard dust (1.1) one has $p=0$, $q^{\mu}=0$, and $\pi^{\mu\nu}=0$ identically, independent of the background or perturbations. In our setup these quantities vanish only on the homogeneous background and, as we will show in the following subsection, are generically nonzero once inhomogeneities are present. Thus the model mimics pressureless dark matter at the background level while allowing nontrivial perturbations.
@@ 4.1 Comoving frame
The acceleration in the comoving frame is given by
$a_{\mu}\equiv u^{\alpha}\nabla_{\alpha}u_{\mu}=-\frac{1}{2X}V_{\mu}\,$, (4.13)
where $V_{\mu}$ is the transverse vector defined in (3.8). Squaring gives
$a_{\mu}a^{\mu}$ $=\frac{1}{4X^{2}}\,V_{\mu}V^{\mu}=\frac{X}{4}\,{\cal C}.$ (4.14)
Note that the above relation is fully covariant and holds in comoving frame. For a generic choice of $u^{\mu}$ not proportional to $\phi^{\mu}$, there is no universal identity relating $a_{\mu}a^{\mu}$ to $\cal C$}.
@@
$\varphi_{\perp}=-\sqrt{-X}\,,\qquad\varphi^{\mu}=0\,,\qquad\chi_{\perp}=\tfrac{Y}{\sqrt{-X}}\,,\qquad\chi^{\mu}=V^{\mu}=-2Xa^{\mu}\,$, (4.15)
@@
$\rho$ $=2\lambda\!\left(C_{\phi\phi}X+C_{\phi X}Y+C_{XX}\tfrac{Y^{2}}{X}\right)\,,\qquad q^{\mu}=2\lambda\,X\sqrt{-X}\left(C_{\phi X}+2C_{XX}\tfrac{Y}{X}\right)a^{\mu}\,$, (4.16)
$\tau^{\mu\nu}$ $=-8\lambda\,C_{XX}\,X^{2}\,a^{\mu}a^{\nu},\quad p=-\tfrac{8\lambda}{3}\,C_{XX}\,X^{2}\,a_{\alpha}a^{\alpha},\quad\pi^{\mu\nu}=-8\lambda\,C_{XX}\,X^{2}\left(a^{\mu}a^{\nu}-\tfrac{1}{3}a_{\alpha}a^{\alpha}\,h^{\mu\nu}\right).$
This frame makes the higher-derivative effects manifest: since $\varphi^{\mu}=0$, any nonzero pressure, energy flux, and anisotropic stress arise entirely from $a^{\mu}$, the spatial projection of $X^{\mu}$. Note that $X_{\mu}=2\,\phi^{\alpha}\phi_{\alpha\mu}$ and $a^{\mu}$ characterizes the spatial second derivatives of the scalar field.
Looking at the subset where the higher-derivative function $f$ is only function of $\cal C$}, defined in Eq. (3.7), the energy-momentum coefficients simplify as shown in Eqs. (3.11) and (3.13) and the hydrodynamical quantities take the following simple forms
$\begin{split}\rho&=2\lambda X\,\widehat{C}_{\phi\phi},\qquad q^{\mu}=0,\qquad p=-\frac{8\lambda}{3}\,f_{\cal C}\,a_{\alpha}a^{\alpha},\\[3.0pt] \pi^{\mu\nu}&=-8\lambda f_{\cal C}\left(a^{\mu}a^{\nu}-\frac{1}{3}a_{\alpha}a^{\alpha}\,h^{\mu\nu}\right)\,.\end{split}$ (4.17)
We see that the energy flux vanishes in this case. This happens because, in the restricted subclass considered here, the higher-derivative corrections are arranged to probe only spatial directions (indeed, $a^{\mu}$ encodes only the spatial higher-derivative structure). This simplification does not hold in the general theory, where a nonzero energy flux is typically present. Therefore, for this subclass the comoving frame coincides with the Landau (energy) frame, defined by $q^{\mu}$=0, which we discuss in the next subsection.
Having derived the results (4.16) and (4.17) in the comoving frame, one can already infer a useful qualitative consequence for linear cosmological perturbations. The key point is that the acceleration $a_{\mu}\equiv u^{\alpha}\nabla_{\alpha}u_{\mu}$ is purely spatial, $u^{\mu}a_{\mu}=0$, and therefore must vanish on an exact FLRW background by homogeneity and isotropy. This is particularly transparent in unitary gauge $\phi=t$, where $a_{\mu}$ has only spatial components ($a_{\mu}\propto\delta_{\mu}^{i}\partial_{i}N$) and hence vanishes at the background level. For the subclass Eq. (4.17), all imperfect pieces induced by the higher-derivative sector are at least quadratic in $a_{\mu}$: the effective pressure and anisotropic stress satisfy $p\propto a_{\alpha}a^{\alpha}$ and $\pi_{\mu\nu}\propto a_{\mu}a_{\nu}-\frac{1}{3}(a_{\alpha}a^{\alpha})h_{\mu\nu}$ (and in this subclass there is no heat flux). It follows that $\delta p$ and $\delta\pi_{\mu\nu}$ vanish at linear order around FLRW, so the higher-derivative corrections do not affect the stress tensor at the level of linear perturbations; deviations from dust-like behavior arise only beyond linear order (while the energy density may still receive corrections through the modified constraint).